# MedUnifier: Unifying Vision-and-Language Pre-training on Medical Data with Vision Generation Task using Discrete Visual Representations

### v2

@@ Abstract
Despite significant progress in Vision-Language Pre-training (VLP), current approaches predominantly emphasize feature extraction and cross-modal comprehension, with limited attention to generating or transforming visual content. This gap hinders the model’s ability to synthesize coherent and novel visual representations from textual prompts, thereby reducing the effectiveness of multi-modal learning. In this work, we propose MedUnifier, a unified VLP framework tailored for medical data. MedUnifier seamlessly integrates text-grounded image generation capabilities with multi-modal learning strategies, including image-text contrastive alignment, image-text matching and image-grounded text generation. Unlike traditional methods that reply on continuous visual representations, our approach employs visual vector quantization, which not only facilitates a more cohesive learning strategy for cross-modal understanding but also enhances multi-modal generation quality by effectively leveraging discrete representations. Our framework’s effectiveness is evidenced by the experiments on established benchmarks, including uni-modal tasks (supervised fine-tuning), cross-modal tasks (image-text retrieval and zero-shot image classification), and multi-modal tasks (medical report generation, image synthesis), where it achieves state-of-the-art performance across various tasks. MedUnifier also offers a highly adaptable tool for a wide range of language and vision tasks in healthcare, marking advancement toward the development of a generalizable AI model for medical applications.
@@ 1 Introduction
The rapid growth of medical imaging datasets has accelerated the development of diverse deep-learning models to enhance clinical decision-making processes, especially for diagnostic radiology. However, annotating these extensive datasets requires specialized expertise, making large-scale annotation economically unfeasible. To overcome this limitation, one effective approach is to leverage information embedded in associated medical reports, which often contain detailed diagnostic descriptions provided by radiologists Reyes et al. (2020). In recent years, deep learning models that utilize multi-modal data as inputs have drawn more attention, driven by advancements in attention mechanisms or transformer-based structures Li et al. (2023d); Zhong et al. (2023); Huh et al. (2023).
@@
Accordingly, numerous cutting-edge vision-and-language pre-training (VLP) models have been developed, many drawing inspiration from the foundational CLIP model Radford et al. (2021). These models primarily leverage a dual-encoder approach, consisting of an image encoder and a text encoder, to independently extract uni-modal features. They aim to maximize cosine similarity between paired image-text data through contrastive learning. Researchers have further enhanced these models by incorporating domain-specific knowledge and making targeted adjustments to the original CLIP architecture, resulting in label-efficient adaptations Zhang et al. (2022); Huang et al. (2021); Zhou et al. (2022a); You et al. (2023); Wu et al. (2023); Wang et al. (2022d). In the dual-encoder framework, the predominant learning objective is achieved through a contrastive loss, which aligns paired data while effectively distinguishing unpaired data within the multi-modal feature space. In addition to the dual-encoder framework, fusion-based encoders have attracted considerable attention. These fusion models utilize self-attention or co-attention mechanisms to achieve early integration of visual and textual modalities Li et al. (2021); Bao et al. (2022); Li et al. (2022a). This joint processing enables the learning of multi-modal representations that are crucial for tasks requiring complex multi-modal reasoning, such as medical visual question answering. For fusion models, cross-modal matching with hard sampling strategies is employed to strengthen correlations between matched data. Along the line image-grounded text understanding with the masked language modelling (MLM), originally developed for the BERT model Devlin (2018), is also deployed to enhance multi-model interaction. PTUnifier Chen et al. (2023b), as a representative, focused on multi-modality understanding yet did not possess generative capabilities during pre-training, thus necessitating the use of an additional language decoder and fine-tuning for downstream language generative tasks. Meanwhile, image-grounded text generation with the causal language modelling (CLM), stemmed from Transformer model Vaswani (2017), is often applied to facilitate vision-grounded language generation tasks. However, we observe that current VLP approaches often overlook the generation of visual content, which limits the model’s capacity to produce coherent and novel visual representations based on textual or multi-modal prompts, thus reducing the potential of multi-modal learning. Although recent studies have integrated masked image modelling (MIM) into the VLP framework Zhou et al. (2022b); Chen et al. (2023a), this approach does not fully enable the generation of comprehensive visual content nor capture detailed visual information.
@@
In this paper, we present MedUnifier, a unified VLP framework for medical data (Figure 1), designed to seamlessly integrate text-grounded image generation with advanced multi-modal learning strategies, including image-text contrastive alignment, image-text matching, and image-grounded text generation. Our approach begins with a Transformer model featuring learnable embeddings, inspired by BLIP-2 Li et al. (2023a), a 12-layer transformer encoder as the trainable component, paired with a pre-trained visual encoder that embeds preliminary visual features. To extend its functionality for vision generation, we introduce a novel learning objective, termed text-grounded image generation (TIG) loss. This objective leverages vector-quantization to facilitate discrete visual representation learning Van Den Oord et al. (2017), guiding vision generation using textual data. Additionally, we devise a novel latent adapter to connect the base model with the image generation module, enabling end-to-end co-training with three other learning objectives: image-text contrastive (ITC), image-text matching (ITM), and image-grounded text generation (ITG) losses. To our knowledge, MedUnifier is the first model to adapt learnable embeddings to the medical domain, bridging the gap between existing VLP paradigms and text-grounded image generation to enhance multi-modal alignment. The main contribution of this study is listed below:
@@
  • We leverage information embedded in textual inputs to support discrete visual representation learning with a bridging design to guide the generation of visual outputs and enhance the interpretation of visual and textual data.
@@
  • We designed special TIG module to capture fine-grained details through recovering pixel-level information from hierarchical multi-modal representations, enabling the model to identify subtle visual details, commonly available in medical data (e.g. small nodules, slight opacities, etc).
@@ 2.1 Vision-and-language Pre-training (VLP)
Vision-language models (VLMs) have attracted considerable attention due to their powerful ability to integrate visual and textual data, significantly enhancing image captioning, visual question answering, and cross-modal retrieval. The predominant paradigms for VLP can be broadly classified into two main categories. The first one focuses on learning uni-modal encoders for text and images Jia et al. (2021); Zhang et al. (2022); Huang et al. (2021); Zhou et al. (2022a); You et al. (2023); Wu et al. (2023); Wang et al. (2022d), respectively. However, this dual-encoder architecture limits the capacity to establish intricate interactions between text and image. Another line of work predominantly focuses on fusion encoder-based structure Wang et al. (2022b); Bao et al. (2022); Su et al. (2019); Li et al. (2021). This method seeks to develop a unified encoder that effectively processes both visual and textual data by leveraging feature fusion strategies to facilitate meaningful interactions between the two modalities. Additionally, In the medical setting, Chen Chen et al. (2023b) proposed an effective framework to unify dual-encode style and fusion-encoder style by constructing a soft prompts pool as placeholders of missing modality. However, these methods do not take the generation of visual information into consideration, and lack exploration of detailed vision content. In this study, we adopt and extend a fusion encoder-based framework to better align visual and textual features by incorporating vector quantization to enable the learning of discrete visual representations, thus facilitating effective vision generation guided by pertinent textual information. Compared to existing studies, our work aligns various modalities and simultaneously creates generic and versatile representations by leveraging the complementary strengths of various losses synergistically, therefore alleviating the additional pre-training stages for expert image tokenizer and iterative denoising.
@@ 2.2 Text to Image (T2I) Generation
Text-to-image (T2I) tasks aim to generate an image according to a given textual description. In the initial stage, generative adversarial networks (GANs) Kang et al. (2023); Li et al. (2022b); Reed et al. (2016); Zhang et al. (2021) and auto-regressive (AR) transformers Ding et al. (2022); Gafni et al. (2022); Ramesh et al. (2021); Yu et al. (2022) were widely recognized for their exceptional performance and popularity. Advancements in variational auto-encoders (VAEs) and vector quantized VAEs (VQ-VAEs) have further improved T2I generation by introducing a discrete latent space for more stable generation Gu et al. (2022); Ramesh et al. (2021); Yu et al. (2022); Ding et al. (2021); Nichol et al. (2021). A LLM-based method, called LLM-CXR Lee et al. (2023b), incorporated pre-trained VQ-GAN into powerful LLM, predicting dual-modal token in joint encoding space. Moreover, diffusion models Dhariwal and Nichol (2021); Ho et al. (2020) have recently taken the leading position in T2I generation tasks. These models harness image space Balaji et al. (2022); Ho et al. (2022); Nichol et al. (2021); Wang et al. (2022c) or latent space Gu et al. (2022); Ramesh et al. (2022); Rombach et al. (2022) for image denoising processes, achieving exceptional image generation quality. By adapting advanced diffusion model into medical domain, MedM2G Zhan et al. (2024) emphasized high-quality content generation by multi-stage diffusion processes. Despite their effectiveness, diffusion models are notably resource-intensive, often necessitating hundreds or thousands of iterative steps for denoising, leading to significantly slower generation speeds. It is hard for diffusion models to get evident visual features due to modelling data distribution implicitly Song et al. (2020). Consequently, for our T2I tasks, we turn to VQ-VAEs to learn more robust representations, thereby enhancing the quality and efficiency of medical image generation effectively.
@@ 3.1 Problem definition
We formulate the Med-VLP problem with inspiration from the previous studies Chen et al. (2023b, 2020). Formally, given a set of medical images $X^{I}\in\left\{x^{i}_{1},x^{i}_{2},\dots,x^{i}_{n}\right\}$ with corresponding clinical reports $X^{T}\in\left\{x^{t}_{1},x^{t}_{2},\dots,x^{t}_{n}\right\}$, the entire pre-training objective function can be defined as
@@
where $\mathcal{F}$ represents backbone that takes the paired $[x^{i},x^{t}]$ as input. $\mathcal{H}_{m}$ stands for task-specific modules for further encoding visual and textual features. $\mathcal{L}_{m}$ and $\lambda_{m}$ are different loss functions and their weights for the overall loss calculation with the total number of loss functions being $M$. The primary goal is to minimize the overall loss $\mathcal{L}_{\text{total}}$ and develop a highly effective parameterized model capable of excelling across a wide range of downstream tasks.
@@ 3.2 Model architecture
We adopted BERT-styled Transformer Li et al. (2023b) as image-text encoder network for fusing multi-modal information in depth. The image-text encoder is constructed by 12 transformer blocks. Its input contains a set of learnable embeddings and clinical reports tokenized by words Devlin (2018). Learnable embeddings and textual input interact with each other through self-attention layers Vaswani (2017) using different masking strategies for various pre-training tasks. We further processed input images into a set of patch embeddings by using a pre-trained, frozen Vision Transformer (ViT) from Fang et al. (2023), attaining informative visual features while saving computational costs during training. Moreover, the initial visual embeddings engage with the image-text encoder network through cross-attention layers, which are integrated within each transformer block. At the final layer of the last transformer block, we incorporated distinct, task-specific heads tailored to address various tasks.
@@
The primary objective of VLP is to learn representations that are transferable across tasks. While the most advanced vision generative models, such as GANs and DMs, excel at generating high-quality and diverse visual content, they typically capture data distributions implicitly. As a result, these models make it challenging to explicitly access intermediate visual representations, limiting their utility in VLP applications. Furthermore medical data often display similar visual or textual patterns and hence encompass nuanced yet crucial medical pattern, which is often overlooked by the existing VLM approaches. To address these constraints, we integrated a vector-quantized variational auto-encoder (VQ-VAE) Van Den Oord et al. (2017) within a cross-modal interactive fusion framework to generate high-quality synthetic visual content and improve performance across multiple downstream tasks.
@@ 3.3 Fusion of visual and textual features
Given that $x_{i}$ out of a paired input $[x^{i},x^{t}]$, $x^{i}\in\mathbb{R}^{C\times H\times W}$, we divided the entire image into $L_{v}$ patches with spatial size $(h,w)$ followed by convolutional operation and added learnable positional encodings:
@@
where $\boldsymbol{p}\in\mathbb{R}^{d_{v}}$ stands for input patch embedding, $\boldsymbol{E}^{v}_{pos}\in\mathbb{R}^{(1+L_{v})\times d_{v}}$ is learnable positional encodings and $L_{v}=\frac{H}{h}*\frac{W}{w}$. We prepended $\boldsymbol{p}_{[CLS]}$ for aggregation of visual information. Then these patch embeddings get passed through a standard pre-trained ViT-g, denoted as $E_{I}$, to attain preliminary visual embeddings $\boldsymbol{f}^{v}\in\mathbb{R}^{(L_{v}+1)\times d_{v}}$:
@@
where $\boldsymbol{f}^{v}_{[CLS]}$ is global visual feature, outputs of patch embeddings $\boldsymbol{f}^{v}_{local}\in\mathbb{R}^{L_{v}\times d_{v}}$ represent local visual features.
For the corresponding textual input, we followed BERT Devlin (2018) to tokenize the input text to word token embeddings, adding learnable positional encodings:
@@
where $\boldsymbol{w}\in\mathbb{R}^{d_{t}}$ represents word token embeddings, $\boldsymbol{E}^{t}_{pos}\in\mathbb{R}^{(1+L_{t})\times d_{t}}$ is learnable positional encodings. We prepended a special token $\boldsymbol{w}_{[SPE]}$ at the beginning and utilized different $[SPE]$ tokens for identifying differing tasks.
In order to enable the interaction between word token embeddings and preliminary visual embeddings, we constructed a set of learnable embeddings, denoted as $\boldsymbol{Q}=[\boldsymbol{q}_{1},\boldsymbol{q}_{2},\dots,\boldsymbol{q}_{L_{q}}],\boldsymbol{Q}\in\mathbb{R}^{L_{q}\times d_{q}}$. We unified word token embeddings and learnable embeddings of the same feature dimension, e.g. $d_{t}=d_{q}$. Then we concatenated $\boldsymbol{Q}$ and $\boldsymbol{X}^{t}$ to form the input of the image-text encoder, denoted as $E_{Q}$, encoding it to get output embeddings:
@@
where $\boldsymbol{f}^{q}\in\mathbb{R}^{L_{q}\times d_{q}}$ is learned embeddings, $\boldsymbol{f}^{t}_{[SPE]}\in\mathbb{R}^{d_{t}}$, $\boldsymbol{f}^{t}_{local}\in\mathbb{R}^{L_{t}\times d_{t}}$ represent special text representation and all word token embeddings, respectively. We employed a cross-modal cross-attention mechanism Alayrac et al. (2022) to facilitate interaction between learnable embeddings and preliminary visual embeddings. This design encourages the learnable embeddings to capture relevant visual information and enables $\boldsymbol{f}^{q}$ to function as the final visual representation. To ensure clarity over the fusion of learnable embeddings and textual representations, we implemented distinct masking strategies within the self-attention layers (Figure 2, right panel).
@@
where $\mathcal{H}_{itc}^{q},\mathcal{H}_{itc}^{t}$ are ITC heads that convert visual and textual representations into joint space. We computed the pairwise similarity between each visual and textual representation $\boldsymbol{g}^{q}$ and $\boldsymbol{g}^{t}$ and chose the highest one as the image-text similarity to calculate bi-directional contrastive loss:
@@
We here expanded visual representation space from a conventional single vector to a set of vectors e.g. $\boldsymbol{f}^{q}\in\mathbb{R}^{L_{q}\times d_{q}}$ which is different from Huang et al. (2021); Wu et al. (2023).
@@
This task aims to learn a precise alignment between visual and textual representations by training a model to classify image-text pairs as either positive (matched) or negative (unmatched) in a binary classification framework. We implemented a bi-directional mask (Figure 2, right panel) that allows all learnable embeddings and word token embeddings to attend to one another. The resulting output of learned embeddings denoted as $\boldsymbol{f}^{q}$, capture enriched multi-modal information. These tokens were then fed into a two-class linear classifier, $\mathcal{H}_{itm}$, where the outputs are averaged across learned embeddings to generate a logit and compute the Image-Text Matching (ITM) loss:
@@
$Y$ represents ground truth labels within mini-batch by hard negative samples mining, as stated in Li et al. (2021).
@@
This task is to train the model to generate text conditioned on paired images. To achieve a coherent and precise generation of medical reports within a unified VLP framework, we chose CLM Radford (2018); Brown (2020) where each word token attends only to preceding tokens, following a GPT-style language model architecture Brown (2020). Inspired by UniLM Dong et al. (2019), we implemented a multi-modal causal self-attention mask (Figure 2, right panel) to control vision-language interactions. For task completion, we replaced the special token $\boldsymbol{w}{[SPE]}$ with $\boldsymbol{w}{[DEC]}$ to signal a decoding task. We also introduced a word token prediction head, denoted as $\mathcal{H}_{itg}$, to predict the current word token. This learning objective is formalized as:

### v3

@@ Abstract
Despite significant progress in Vision-Language Pre-training (VLP), current approaches predominantly emphasize feature extraction and cross-modal comprehension, with limited attention to generating or transforming visual content. This gap hinders the model’s ability to synthesize coherent and novel visual representations from textual prompts, thereby reducing the effectiveness of multi-modal learning. In this work, we propose MedUnifier, a unified VLP framework tailored for medical data. MedUnifier seamlessly integrates text-grounded image generation capabilities with multi-modal learning strategies, including image-text contrastive alignment, image-text matching and image-grounded text generation. Unlike traditional methods that reply on continuous visual representations, our approach employs visual vector quantization, which not only facilitates a more cohesive learning strategy for cross-modal understanding but also enhances multi-modal generation quality by effectively leveraging discrete representations. Our framework’s effectiveness is evidenced by the experiments on established benchmarks, including uni-modal tasks, cross-modal tasks, and multi-modal tasks, where it achieves state-of-the-art performance across various tasks. MedUnifier also offers a highly adaptable tool for a wide range of language and vision tasks in healthcare, marking advancement toward the development of a generalizable AI model for medical applications.
@@ 1 Introduction
The rapid growth of medical imaging datasets has accelerated the development of deep-learning models to enhance clinical decision-making processes. However, annotating these extensive datasets requires specialized expertise, making large-scale annotation unfeasible. To overcome this limitation, one effective approach is to leverage associated medical reports containing detailed diagnostic descriptions provided by radiologists [52]. In recent years, deep learning models that utilize multi-modal data as inputs have drawn more attention, driven by advancements in attention mechanisms or transformer-based models [40, 70, 27].
@@
Accordingly, vision-and-language pre-training (VLP) models have been developed, many drawing inspiration from the foundational CLIP model [48]. These models primarily leverage a dual-encoder approach, consisting of an image encoder and a text encoder, to independently extract uni-modal features. They aim to maximize cosine similarity between paired data via contrastive learning. Researchers have further enhanced these models by incorporating domain-specific knowledge and making targeted adjustments to the original CLIP, resulting in label-efficient adaptations [69, 26, 71, 64, 63, 62]. In addition, fusion-based encoders have attracted considerable attention. These fusion models utilize self-attention or co-attention mechanisms to achieve early integration of visual and textual modalities [34, 3, 35]. This joint processing enables the learning of multi-modal representations that are crucial for tasks requiring complex multi-modal reasoning, such as medical visual question answering. For fusion models, cross-modal matching with hard sampling strategies is employed to strengthen correlations between matched data. Image-grounded text understanding with masked language modelling (MLM), originally developed for the BERT [13], is also deployed to enhance multi-model interaction. PTUnifier [10], as an example, focused on multi-modal understanding yet did not possess generative ability during pre-training, thus necessitating the use of an additional language decoder and fine-tuning for language generative tasks. Meanwhile, image-grounded text generation with the causal language modelling (CLM) is often applied to facilitate vision-grounded language generation tasks. However, we observe that current VLP approaches often overlook the generation of visual content, limiting the model’s capacity to produce coherent and novel visual representations based on textual or multi-modal prompts, thus reducing the potential of multi-modal learning. Although recent studies have integrated masked image modelling (MIM) into the VLP framework [72, 7], this approach does not fully enable the generation of comprehensive visual content nor capture detailed visual information effectively.
@@
In this paper, we present MedUnifier, a unified VLP framework for medical data (Figure 1), designed to seamlessly integrate text-grounded image generation with advanced multi-modal learning strategies, including image-text contrastive alignment, image-text matching, and image-grounded text generation. Our approach begins with a Transformer model featuring learnable embeddings, inspired by BLIP-2 [36], a 12-layer transformer encoder as the trainable component, paired with a pre-trained visual encoder that embeds preliminary visual features. To extend its functionality for vision generation, we introduce a novel learning objective, termed text-grounded image generation (TIG) loss. This objective leverages vector-quantization to facilitate discrete visual representation learning [56], guiding vision generation using textual data. Additionally, we devise a novel latent adapter to connect the base model with the image generation module, enabling end-to-end co-training with three other learning objectives: image-text contrastive (ITC), image-text matching (ITM), and image-grounded text generation (ITG) losses. To our knowledge, MedUnifier is the first model to adapt learnable embeddings to the medical domain, bridging the gap between existing VLP paradigms and text-grounded image generation to enhance multi-modal alignment. The main contribution of this study is listed below:
@@
  • We designed a novel TIG module to capture fine-grained details by recovering pixel-level information from hierarchical multi-modal representations, enabling the model to identify subtle visual details, commonly available in medical data (e.g. small nodules, slight opacities, etc).
@@ 2.1 Vision-and-language Pre-training (VLP)
Vision-language models (VLMs) have attracted considerable attention due to their powerful ability to integrate visual and textual data, significantly enhancing image captioning, visual question answering, and cross-modal retrieval. The predominant paradigms for VLP can be broadly classified into two main categories. The first one focuses on learning uni-modal encoders for text and images [29, 69, 26, 71, 64, 63, 62], respectively. However, this dual-encoder architecture limits the capacity to establish intricate interactions between text and image. Another line of work predominantly focuses on fusion encoder-based structure [60, 3, 55, 34] to facilitate meaningful interactions between the two modalities. In the medical setting, Chen et al. [10] proposed an effective framework to unify dual-encode style and fusion-encoder. However, these methods do not take the generation of visual information into consideration, and lack exploration of detailed vision content. In this study, we adopt and extend a fusion encoder-based framework to better align visual and textual features by incorporating vector quantization to enable the learning of discrete visual representations, thus facilitating effective vision generation guided by pertinent textual information. Compared to existing studies, our work aligns various modalities and simultaneously creates generic and versatile representations by leveraging the complementary strengths of various losses synergistically, therefore alleviating the additional pre-training stages for expert image tokenizer and iterative denoising.
@@ 2.2 Text to Image (T2I) Generation
Text-to-image (T2I) tasks aim to generate an image according to a given textual description. Generative adversarial networks (GANs) [31, 39, 51, 68] and auto-regressive (AR) transformers [16, 19, 49, 65] were widely recognized for their exceptional performance and popularity. Advancements in variational auto-encoders (VAEs) and vector quantized VAEs (VQ-VAEs) have further improved T2I generation by introducing a discrete latent space for more stable generation [20, 49, 65, 15, 45]. LLM-CXR [33] incorporated pre-trained VQ-GAN into powerful LLM, predicting dual-modal tokens. Moreover, diffusion models [14, 25] have recently taken the leading position in T2I generation tasks. By adapting the advanced diffusion model, MedM2G [67] emphasized high-quality content generation. Despite their effectiveness, diffusion models are notably resource-intensive, often necessitating thousands of iterative steps for denoising, leading to significantly slower speeds. It is hard for diffusion models to get evident visual features due to modelling data distribution implicitly [54]. Consequently, for our T2I tasks, we turn to VQ-VAEs to learn more robust representations, thereby enhancing the quality and efficiency of medical image generation effectively.
@@ 3.1 Problem definition
We formulate the Med-VLP problem with inspiration from the previous studies [10, 9]. Formally, given a set of medical images $X^{I}\in\left\{x^{i}_{1},x^{i}_{2},\dots,x^{i}_{n}\right\}$ with corresponding clinical reports $X^{T}\in\left\{x^{t}_{1},x^{t}_{2},\dots,x^{t}_{n}\right\}$, the entire pre-training objective function can be defined as
@@
where $\mathcal{F}$ represents backbone taking the paired $[x^{i},x^{t}]$ as input. $\mathcal{H}_{m}$ stands for task-specific modules for further encoding visual and textual features. $\mathcal{L}_{m}$ and $\lambda_{m}$ are different loss functions and their weights for the overall loss calculation with the total number of loss functions being $M$.
@@ 3.2 Model architecture
We adopted BERT-styled Transformer [37] as image-text encoder network for fusing multi-modal information in depth. The image-text encoder is constructed by 12 transformer blocks. Its input contains a set of learnable embeddings and clinical reports tokenized by words [13]. Learnable embeddings and textual input interact with each other through self-attention layers [57] using different masking strategies for various pre-training tasks. We further processed input images into a set of patch embeddings by using a pre-trained, frozen Vision Transformer (ViT) from [18], attaining informative visual features while saving computational costs during training. Moreover, the initial visual embeddings engage with the image-text encoder network through cross-attention layers, which are integrated within each transformer block. At the final layer of the last transformer block, we incorporated distinct, task-specific heads tailored to address various tasks.
@@
The primary objective of VLP is to learn representations that are transferable across tasks. While the most advanced vision generative models, such as GANs and DMs, excel at generating high-quality and diverse visual content, they typically capture data distributions implicitly, which these models make it challenging to explicitly access intermediate visual representations, limiting their utility in VLP applications. Furthermore medical data often display similar visual or textual patterns and hence encompass nuanced yet crucial medical pattern, which is often overlooked by the existing VLM approaches. To address these constraints, we integrated a vector-quantized variational auto-encoder (VQ-VAE) [56] within a cross-modal interactive fusion framework to generate high-quality synthetic visual content and improve performance across multiple downstream tasks.
@@ 3.3 Fusion of visual and textual features
Given that $x^{i}\in\mathbb{R}^{C\times H\times W}$, we divided the entire image into $L_{v}$ patches with spatial size $(h,w)$ through convolutional operation and added learnable positional encodings:
@@
We prepended $\boldsymbol{p}_{[CLS]}$ for aggregation of visual information. Then these patch embeddings get passed through a standard pre-trained ViT-g, denoted as $E_{I}$, to attain preliminary visual embeddings $\boldsymbol{f}^{v}\in\mathbb{R}^{(L_{v}+1)\times d_{v}}$:
@@
where $\boldsymbol{f}^{v}_{[CLS]}$ is global visual feature, $\boldsymbol{f}^{v}_{local}\in\mathbb{R}^{L_{v}\times d_{v}}$ represent local visual features.
For the textual input, we followed BERT [13] to tokenize the input text to word embeddings, adding learnable positional encodings:
@@
We prepended a special token $\boldsymbol{w}_{[SPE]}$, utilizing different $[SPE]$ tokens for identifying differing tasks.
In order to enable the interaction between word embeddings and preliminary visual embeddings, we constructed a set of learnable embeddings, denoted as $\boldsymbol{Q}=[\boldsymbol{q}_{1},\boldsymbol{q}_{2},\dots,\boldsymbol{q}_{L_{q}}],\boldsymbol{Q}\in\mathbb{R}^{L_{q}\times d_{q}}$. We unified word embeddings and learnable embeddings of the same feature dimension, e.g. $d_{t}=d_{q}$. Then we concatenated $\boldsymbol{Q}$ and $\boldsymbol{X}^{t}$ to form the input of the image-text encoder, denoted as $E_{Q}$, encoding it to get output embeddings:
@@
We employed a cross-attention mechanism [1] to facilitate interaction between learnable embeddings and preliminary visual embeddings. This design enables $\boldsymbol{f}^{q}$ to function as the final visual representation. To ensure clarity over the fusion of learnable embeddings and textual representations, we implemented distinct masking strategies within the self-attention layers (Figure 2, right panel).
@@
where $\mathcal{H}_{itc}^{q},\mathcal{H}_{itc}^{t}$ are ITC heads. We computed the pairwise similarity between each visual and textual representation $\boldsymbol{g}^{q}$ and $\boldsymbol{g}^{t}$ and chose the highest one as the image-text similarity to calculate bi-directional contrastive loss:
@@
We here expanded visual representation space from a conventional single vector to a set of vectors e.g. $\boldsymbol{f}^{q}\in\mathbb{R}^{L_{q}\times d_{q}}$ which is different from [26, 63].
@@
This task aims to learn a precise alignment by classifying image-text pairs as either positive or negative. We implemented a bi-directional mask (Figure 2, right panel) that allows all learnable embeddings and word token embeddings to attend to one another. The resulting output of learned embeddings denoted as $\boldsymbol{f}^{q}$, capture enriched multi-modal information. These tokens were then fed into a two-class linear classifier, $\mathcal{H}_{itm}$, where the outputs are averaged across learned embeddings to generate a logit and compute the Image-Text Matching (ITM) loss:
@@
$Y$ represents ground truth labels within mini-batch by hard negative samples mining, as stated in [34].
@@
This task is to generate text conditioned on paired images. To achieve a coherent and precise generation of medical reports within a unified VLP framework, we chose CLM [47, 5] where each word token attends only to preceding tokens, following a GPT-style language model architecture [5]. Inspired by UniLM [17], we implemented a multi-modal causal self-attention mask (Figure 2, right panel). We replaced the special token $\boldsymbol{w}{[SPE]}$ with $\boldsymbol{w}{[DEC]}$ to signal a decoding task. We also introduced a word prediction head, denoted as $\mathcal{H}_{itg}$. This learning objective is formalized as: